OpenStudy (anonymous):

solve: 3^4x-7 = 4^2x+3

OpenStudy (anonymous):

81x-7 = 16x +3

OpenStudy (anonymous):

65x=10

OpenStudy (anonymous):

x=10/65 and simplify x=2/13

OpenStudy (earthcitizen):

i think 3^(4x-7)=4^(2x+3)

OpenStudy (anonymous):

yeah that's what I meant, I'm new to the website so I didn't know how to input it properly

OpenStudy (earthcitizen):

take logs of both sides, (4x-7)log3=2(2x+3)log2 =4xlog3-7log3=4xlog2+6log2 collect like terms, 4xlog3-4xlog2=6log2+7log3 4x(log3-log2)=6log2+7log3 therefore, x=6log2+7log3/4(log3-log2)

OpenStudy (earthcitizen):

x=(6log2+7log3)/(4(log3-log2))

OpenStudy (anonymous):

ooops..